OpenStudy (anonymous):

What is the vertex form of y= 4x^2+20x+17? And how would you do it? (Please provide steps)

OpenStudy (anonymous):

-5/2,-8

OpenStudy (anonymous):

What would you get for the vertex form though? y= a(x+h)^2+k

OpenStudy (amistre64):

parse out the x^2+x part

OpenStudy (amistre64):

y= 4x^2 +20x +17 y= (4x^2 +20x) +17 ; now divide out the 4 y= 4(x^2 +5x) +17; and complete the square

OpenStudy (amistre64):

y= 4(x^2 +5x +(5/2)^2) +17-4(5/2^2) y= 4(x^2 +5x +(5/2)^2) +17-4(25/4) y= 4(x^2 +5x +(5/2)^2) +17-25; and compact the square y= 4(x+(5/2))^2 -8

OpenStudy (anonymous):

Thanks!

OpenStudy (amistre64):

yep